Question 14 of 14 g feaming The robot HooRU is lost in space, floating around aimlessly, and radiates heat into the depths of the cosmos at the rate of 14.9 W. HooRUS surface area is 1.65 m2 and the emissivity of its surface is 0.259. Ignore the radiation HooRU absorbs from the cold universe. What is HooRUs temperature? Number
Answer #1

power emitted = P = sigme*e*A*T^4


sigma = stefans constant = 5.67*10^-8w/m^2-K^4

e = 0.259


P = 14.9 W

A = 1.65 m^2

T^4 = P/sigma*e*A

T^4 = 14.9/(5.67*10^-8*0.259*1.65)

T = 157.47 K <------answer